# Emergent vorticity asymmetry of one and two-layer shallow water system captured by a next-order balanced model

### v1

@@ Abstract
The quasigeostrophic (QG) system has served as a useful simplified model for understanding geophysical physical fluid phenomena. The fact that it is based on only one prognostic variable — potential vorticity (PV) — is a simplification that has facilitated much theoretical understanding. However, although QG captures many geophysical turbulence phenomena, it misses important features that occur in shallow water systems at finite Rossby numbers, even those that are “balanced”. For example, QG does not capture the emergent vorticity asymmetry or the finite divergence of the velocity fields. Here we present a next-order-in-Rossby extension of QG in the single-layer and multi-layer shallow water context: SWQG+1. A freely decaying simulation shows that SWQG+1 can capture the negatively skewed vorticity found in simulations of the shallow water model in the same setting. We also extend the model to the multi-layer configuration. Simulation of nonlinear evolution of a baroclinically unstable jet in SWQG+1 shows that it can capture vorticity asymmetry and finite divergence of strain-driven fronts. We propose here that SWQG+1 could be a useful intermediate model for theoretical and numerical study of the balanced dynamics of geophysical physical fluid dynamics, with possible applications for forward and inverse modeling of atmospheres and oceans on Earth and other planets.
@@ 1 Introduction
The shallow water system has been a crucial simple model for our understanding of geophysical fluid dynamics (GFD) (e.g. Zeitlin, 2018). It simplifies the more complex GFD equations (e.g. the Boussinesq system) by assuming that one can represent the vertical variation using only one (or a few) stacked stratified layers. This simplification allows for easier numerical and theoretical investigations of its properties, and in many cases shallow water captures the realistic properties of geophysical phenomena in the atmosphere and ocean of earth and other planets (e.g. Dowling & Ingersoll, 1989; Polvani et al., 1994; Cho & Polvani, 1996; Lambaerts et al., 2012; Bembenek et al., 2020, for just a few).
@@
A further simplification of the shallow water system is the shallow water quasi-geostrophic equation (SWQG) (Vallis, 2017; Zeitlin, 2018). It assumes rotation dominates and thus the geostrophic balance is valid and the Rossby number is small. SWQG is then an asymptotic approximation of the shallow water model in the limit of vanishing Rossby number. It is a prototypical example of a balanced model, where it can be written with only a single prognostic variable, potential vorticity (PV), from which all other physical variables can be diagnosed. This has the consequence that gravity waves are not permitted in its dynamics. However, SWQG (and its special case, the two-dimensional Euler equation) still has rich turbulence phenomena that closely resemble the balanced dynamics of rotationally dominated atmosphere and ocean. Therefore, they have been the main playground for the study of geophysical turbulence (e.g. Gill et al., 1974; Salmon, 1980; Maltrud & Vallis, 1991; Smith et al., 2002; Dritschel & McIntyre, 2008; Early et al., 2011). The cumulated study of all aspects of SWQG has lead to progress in many practical geophysical applications including the parametrization of eddies (Held & Larichev, 1996; Srinivasan & Young, 2014; Grooms et al., 2015; Gallet & Ferrari, 2020, 2021), the interpolation and interpretation of satellite altimetry measurement of sea surface height (Stammer, 1997; Chelton et al., 2007; Ubelmann et al., 2015), analysis of ocean coherent vortices over topography features (Bretherton & Haidvogel, 1976; Carnevale & Frederiksen, 1987; Siegelman & Young, 2023; LaCasce et al., 2024), the statistics of atmospheric blocking (Marshall & Molteni, 1993; Lucarini & Gritsun, 2020), and the mechanics of jets and spots on gas giants (Turkington et al., 2001; Majda & Wang, 2006; Siegelman et al., 2022; Pizzo & Salmon, 2024).
@@
Despite the vast amount of phenomenology SWQG can capture, it still misses many features of the full shallow water model, even those that might be categorized as “balanced”. This paper focuses on one of such deficiencies: SWQG misses the asymmetry between cyclonic and anticyclonic vortices. Polvani et al. (1994) note that the vorticity skews negative in a freely decaying simulation of the shallow water system. Arai & Yamagata (1994); Graves et al. (2006) further analyze the detailed free evolution of vortices in the shallow water model and conclude that anti-cyclone is more robust to the perturbation of a single vortex and to the influence of other vortices, providing a plausible mechanics of the negative skewness of vorticity. For an ocean gyre forced by wind, Thiry et al. (2024) found the shallow water model version exhibits asymmetric drift of the strong eastward jet, while the SWQG version lacks this asymmetry. For multi-layer shallow water, Lambaerts et al. (2012) find the evolution of baroclinically unstable jets in two-layer shallow water exhibit complex vorticity asymmetry evolution, further complicated by their inclusion of moisture.
@@
In this paper, we aim at a balanced model of the shallow water system that could capture the evolution of asymmetry between cyclonic and anticyclonic vortices in the same parameter space as SWQG. As a by-product, we also wish to capture the finite divergence of shallow water flows, missed by SWQG. Some balanced models already exist in the literature. One line of models includes the frontal geostrophic model by Cushman-Roisin (1986) and the generalized geostrophic equation by Cushman-Roisin & Tang (1989, 1990). These models capture a more extended parameter space of the shallow water model by allowing for large height deviation and can capture vorticity asymmetry. However, they are not higher order in Rossby number than SWQG, and they all insist that the velocities are in geostrophic balance with the height field and thus do not allow for divergent flow. Another model is the Balance Equation (Allen et al., 1990; Spall & Mcwilliams, 1992). It is a capable model of shallow water in the QG regime that includes higher-order effects and is useful as a way to initialize a shallow water simulation with minimal generation of gravity waves. We use it to generate the initial conditions for our freely decaying simulation. However, evolving the Balance Equation is challenging. The system has two time derivatives but is still a balanced model in the sense that there is only one prognostic variable. This is because the two time evolution equations are in fact not independent from each other but coupled in complex nonlinear ways. The time evolution usually needs iterative solvers per timestep (Barth et al., 1990). This poses challenges when using it to study the evolution of the shallow water model. Last but not least is the semi-geostrophic model of Hoskins (1975) adapted to the shallow water context (Cloke & Cullen, 1994; Roulstone & Sewell, 1996). It has features and deficiencies similar to those of the Boussinesq semi-geostrophic model. It has energy and PV conservation and can capture vorticity asymmetry. However, it is not of higher order in Rossby number. The semi-geostrophic model is less accurate in modeling curved fronts. The kinetic energy of the semi-geostrophic model includes only the geostrophically balanced component, making it hard to interpret. The inversion from PV involves a change of coordinate and solving a nonlinear elliptic Monge-Ampère problem, both are hard to do numerically. Interestingly, the mathematical challenge of analyzing a Monge-Ampère problem coupled to a transport equation has inspired deep mathematical analysis of the semi-geostrophic model for shallow water (see Cullen & Gangbo, 2001; De Philippis & Figalli, 2014, and references therein).
@@
We propose a balanced model that extends the SWQG into one higher order in Rossby number that we call SWQG+1. It gets its name from the Boussinesq version of the QG+1 model by Muraki et al. (1999). However, the one-layer version of our SWQG+1 model is first derived by Warn et al. (1995) and further explored by Vallis (1996, hereafter V96). The equivalence of the Boussinesq version of the QG+1 model of the model in the Boussinesq version of the V96 model is shown in et al. (2024). Recently Chouksey et al. (2023) has studied a variant of the Warn et al. (1995) model’s ability for balanced initialization of the shallow water model. We take crucial inspiration from Muraki et al. (1999)’s version of QG+1 and rewrite the Warn et al. (1995) model using “potentials”. This potential form allows us to extend the model to multi-layer for the first time. The higher-order in Rossby terms models the asymmetric correlation between height deviation and vorticity and the cyclostrophic balance. These nonlinear effects allow QG+1 to capture the asymmetric evolution of vorticity in the one and multi-layer shallow model, as well as flow fields with finite divergence.
@@
SWQG+1, like QG+1 in the Boussinesq context, is a PV-based balanced model, where the prognostic variable is PV and all other variables are diagnosed from it. Warn et al. (1995) first considered abstractly the problem of constructing higher-order balanced models and determined that a stable model should not expand the prognostic variable. For SWQG+1 this means PV is not expanded in the small Rossby number. All other variables are expanded and “slaved” to PV by the principle of “PV inversion” of Hoskins et al. (1985). Choosing PV as the central variable is also based on the fact that PV has been crucial in understanding many features of the shallow water model and SWQG (Bretherton & Haidvogel, 1976; Haynes & McIntyre, 1987; Hoskins, 1991; Dritschel & McIntyre, 2008). It is acknowledged that the Boussinesq QG+1 currently lack an energy conservation principle. The same is the case for the SWQG+1 model.
@@
In this paper, we derive the one and two-layer SWQG+1 model using the potentials. For one layer, this is an alternative form of the model of Warn et al. (1995) and V96. We explore its turbulent free decay, compared to the corresponding shallow water model. It is shown that SWQG+1 captures the emergent negative vorticity skewness in the shallow water model. We also track the energy evolution. It is shown that while theoretical energy conservation law is lacking, the energy of SWQG+1 is well-behaved and models the energy evolution of the shallow water model well. The one-layer study is contained in Section 2. In Section 3, we extend the SWQG+1 model to many layers for the first time. A simulation of a baroclinically unstable jet demonstrates SWQG+1 can capture the vorticity asymmetry in two layers. Contrary to the results of the one-layer simulations, the two-layer simulation shows the initial baroclinic growth stage has a cyclonic bias for vorticity. It also shows patterns of vorticity and divergence similar to strain-driven fronts. The vortex stretching mechanism leads to the cyclonic vorticity bias. We conclude in Section 4 and discuss possible applications and extensions to the model.
@@ 2 Single-layer shallow water
This section re-derive the SWQG+1 model, providing an alternative to the Warn et al. (1995)’s approach. By using “potentials” inspired by the Muraki et al. (1999) model for Boussinesq, our form of SWQG+1 involved three elliptic inversions where the elliptic operator to be inverted are the same as the one in SWQG, the Screened Poisson operator. We simulate the free decay of random balanced initial conditions of SWQG+1 model as well as the shallow water model and compare their turbulent statistics. SWQG+1 can capture the vorticity asymmetry in the shallow water model well.
@@ 2.1.1 The shallow water system and its properties
We start with the one-layer shallow water model with horizontal velocity $(u,v)$ and water layer height perturbation $h$:
@@
To facilitate future asymptotic analysis but also keep the physical constants for interpretability, we write the equation using the notation of V96. All nondimensional constants are enveloped in curly brackets and one can ignore them if one wants only the physical equations. One can instead obtain the nondimensional equations by removing all physical constants, here $f$, $\beta$, $g$, and $H$. The nondimensional numbers in this model are the Rossby number:
@@
To prepare for the derivation for the SWQG+1 model, we propose a form of representing the three variables of shallow water using three “potentials”
@@
This procedure and naming are inspired by QG+1 for the Boussinesq model where the three potentials are the vector potential form of the three dimensional incompressible velocity field (Muraki et al., 1999; Dù et al., 2024). Here for the shallow water case, the natural counterpart is the two-dimensional Helmholtz decomposition as adopted by Warn et al. (1995) and V96. Though our potentials might look unnatural at first glance, it is useful for the derivation of SWQG+1, where it turns out the elliptic inversion problems for all three potentials are the same as SWQG’s Screened Poisson problems. This uniformity guides the derivation of SWQG+1 in the more complex multi-layer case.
@@ 2.1.2 The shallow water QG and QG+1 model
SWQG+1 is a model that makes the same asymptotic assumption of SWQG but extends it to the next level in Rossby number. That is, we assume the small parameter
@@
$\Phi^{0}$ and $\Phi^{1}$ should have zero mean for all time in the simulation to maintain mass conservation (15c) (assuming $\left\langle{h}\right\rangle=0$). Therefore for SWQG+1 the $\Phi^{0}$ inversion (18b) should use $q-\left\langle{q}\right\rangle$. Note that with $\beta=0$, the integral of the right-hand side is proportional to the total energy at the QG level after integration by parts
@@
Together we have the elliptic inversion for $F^{1}$:
@@
In summary, the SWQG+1 inversions provides a way to obtain the physical variables (15) from knowledge of the PV field $q$. It consists of four linear elliptic inversions (19b), (20), (27), and (28). These inversions are all the same Screened Poisson problem. This is typical for a model based on asymptotic expansions. The above derivation of SWQG+1 uses the same ingredient as the model in Warn et al. (1995) and V96 and they are in fact equivalent. We show this explicitly in Appendix A.
SWQG+1 uses PV as the prognostic variable. Therefore it retains the PV conservation of (6). However, we have not been able to show that it conserved any definition of energy. Instead, we will demonstrate numerically that the usual definition of shallow water energy (9) is well-behaved though not conserved. SWQG+1 does not conserve potential enstrophy (7) either, which additionally requires the height evolves following the height equation (1c). SWQG+1’s height is inverted from PV and does not necessarily follow (1c). We also show from simulations that the potential enstrophy (7) behaves within expectation. Overall, even though SWQG+1 does not explicitly conserve the energy and potential enstrophy, they mirror the behavior of the underlying shallow water model. Theoretical thinking based on the conserved quantities can be applied to the SWQG+1 simulations.

### v3

@@ Abstract
The Quasi-Geostrophic (QG) system has served as a useful simplified model for understanding geophysical fluid phenomena. The fact that it is based on only one prognostic variable — potential vorticity (PV) — is a simplification that has facilitated much theoretical understanding. However, although QG captures many geophysical turbulence phenomena, it misses important features that occur in shallow water systems at finite Rossby numbers, even those that are “balanced”. For example, QG does not capture the emergent vorticity asymmetry or the finite divergence of the velocity fields. Here we present a next-order-in-Rossby extension of QG in the single-layer and multi-layer shallow water context: SWQG+1. A freely decaying simulation shows that SWQG+1 can capture the negatively skewed vorticity found in simulations of the shallow water model in the same setting. We also extend the model to the multi-layer configuration. Simulations of nonlinear evolution of a baroclinically unstable jet in SWQG+1 show that it can capture vorticity asymmetry and finite divergence of strain-driven fronts.
@@ 1 Introduction
The shallow water system has been a crucial workhorse model for our understanding of geophysical fluid dynamics (GFD) (e.g. Zeitlin, 2018). It simplifies the more complex GFD equations (e.g., the Boussinesq system) by assuming that one can represent the vertical variation using only one (or a few) stacked stratified layers. This simplification allows for easier numerical and theoretical investigations of its properties, and in many cases shallow water captures the realistic properties of geophysical phenomena in the atmosphere and ocean of Earth and other planets (e.g. Dowling & Ingersoll, 1989; Polvani et al., 1994; Cho & Polvani, 1996; Lambaerts et al., 2012; Bembenek et al., 2020, for just a few).
@@
A further simplification of the shallow water system is the Shallow Water Quasi-Geostrophic equation (SWQG) (Vallis, 2017; Zeitlin, 2018). It assumes rotation dominates and thus the geostrophic balance approximation is valid and the Rossby number is small. SWQG is then an asymptotic approximation of the shallow water model in the limit of vanishing Rossby number. It is a prototypical example of a balanced model, in that it can be written with only a single prognostic variable, potential vorticity (PV), from which all other physical variables can be diagnosed. This has the consequence that gravity waves are not permitted in its dynamics. SWQG (and its special case, the two-dimensional Euler equations) can capture the dynamics of shallow water in surprisingly wide parameter regime (Sugimoto et al., 2007), and it has rich turbulence phenomena that closely resemble the balanced dynamics of rotationally dominated atmospheres and oceans (e.g. Gill et al., 1974; Salmon, 1980; Maltrud & Vallis, 1991; Smith et al., 2002; Dritschel & McIntyre, 2008; Early et al., 2011). The study of SWQG has led to progress in many practical geophysical applications, including the parametrization of eddies (Held & Larichev, 1996; Thompson & Young, 2007; Srinivasan & Young, 2014; Gallet & Ferrari, 2020, 2021), the interpolation and interpretation of satellite altimetry measurement of sea surface height (Stammer, 1997; Chelton et al., 2007; Ubelmann et al., 2015), the analysis of ocean coherent vortices over topography features (Bretherton & Haidvogel, 1976; Carnevale & Frederiksen, 1987; Siegelman & Young, 2023; LaCasce et al., 2024), the statistics of atmospheric blocking (Marshall & Molteni, 1993; Lucarini & Gritsun, 2020), and the mechanics of jets and spots on gas giants (Turkington et al., 2001; Majda & Wang, 2006; Siegelman et al., 2022; Pizzo & Salmon, 2024), to name but a few.
@@
Despite the range of phenomenology SWQG captures, it still misses many features of the full shallow water model, even those that might be categorized as “balanced”. SWQG misses the asymmetry between cyclonic and anticyclonic vortices. Polvani et al. (1994) note that the vorticity skews negative in a freely decaying simulation of the shallow water system. Arai & Yamagata (1994); Graves et al. (2006) further analyze the detailed free evolution of vortices in the shallow water model and conclude that anticyclones are more robust to small-scale perturbation and to the influence of other vortices, providing a plausible mechanics of the negative skewness of vorticity. For an ocean gyre forced by wind, Thiry et al. (2024) found the shallow water model version exhibits asymmetric drift of the strong eastward jet, while the SWQG version lacks this asymmetry. For multi-layer shallow water, Lambaerts et al. (2012) find that the evolution of baroclinically unstable jets in two-layer shallow water exhibits complex vorticity asymmetry evolution, further complicated by their inclusion of moisture.
@@
In this paper, we aim at a balanced model of the shallow water system that could capture the evolution of asymmetry between cyclonic and anticyclonic vortices. We focus on the parameter space where SWQG is formally valid but lacking in captured phenomena. As a by-product, we also wish to capture the finite divergence of shallow water flows, missed by SWQG. Some balanced models already exist in the literature. One line of models includes the frontal geostrophic model by Cushman-Roisin (1986) and the generalized geostrophic equation by Cushman-Roisin & Tang (1989, 1990). These models capture a more extended parameter space of the shallow water model by allowing for large height deviation and can capture vorticity asymmetry. However, they are not higher order in Rossby number than SWQG, and they all insist that the velocities are in geostrophic balance with the height field and thus do not allow for divergent flow. Another model is the Balance Equation (Allen et al., 1990; Spall & Mcwilliams, 1992). It is a capable model of shallow water in the QG regime that includes higher-order effects and is useful as a way to initialize a shallow water simulation with minimal generation of gravity waves. We use it to generate the initial conditions for our freely decaying simulation. However, evolving the Balance Equation is challenging. The system has two time derivatives but is still a balanced model in the sense that there is only one prognostic variable. This is because the two time evolution equations are not independent from each other but coupled in complex nonlinear ways. The time evolution usually needs iterative solvers per timestep (Barth et al., 1990). This poses challenges when using it to study the evolution of the shallow water model and fully turbulent simulations of it are rare. Yet another model is the semi-geostrophic model of Hoskins (1975) adapted to the shallow water context (Cloke & Cullen, 1994; Roulstone & Sewell, 1996). It has features and deficiencies similar to those of the Boussinesq semi-geostrophic model. It has energy and PV conservation and can capture vorticity asymmetry. However, it is not of higher order in Rossby number. The semi-geostrophic model is less accurate in modeling curved fronts. The kinetic energy of the semi-geostrophic model includes only the geostrophically balanced component, making it hard to interpret. The inversion from PV involves a change of coordinate and solving a nonlinear elliptic Monge-Ampère problem, both are hard to do numerically. Practically, Fletcher (2004) explored its utility as a model for data assimilation in numerical weather prediction. Theoretically, the mathematical challenge of analyzing a Monge-Ampère problem coupled to a transport equation has inspired deep mathematical analysis of the semi-geostrophic model for shallow water (see Cullen & Gangbo, 2001; De Philippis & Figalli, 2014, and references therein). Finally, the optimal balance algorithm of Viúdez & Dritschel (2004) deals more broadly with the issue of optimal separation of balanced and unbalanced components of a snapshot of the flow, going beyond asymptotics (Masur & Oliver, 2020; Chouksey et al., 2023). The usage of the near-optimally diagnosed balanced component of the flow in the time evolution of the model has not been explored. Therefore, for now, it does not fit our bill of studying the gravity-wave free evolution of the balanced component of the shallow water system.
@@
We explore and extend a balanced model that extends the SWQG into one higher order in Rossby number that we call SWQG+1. Warn et al. (1995) first derived SWQG+1for one-layer shallow water system, and Vallis (1996, hereafter V96) extended it to include the $\beta$-effect. The higher-order-in-Rossby terms model the asymmetric correlation between height deviation and vorticity and the cyclostrophic balance. Recently Chouksey et al. (2023) has studied a variant of the Warn et al. (1995) model’s ability for balanced initialization of the shallow water model and has connected it to the optimal balance algorithm. In this paper, we rewrite the Warn et al. (1995) model using “potentials”, inspired by Muraki et al. (1999)’s version of QG+1of the Boussinesq system. This new potential form allows us to extend the model to multi-layer for the first time. We show through nonlinear simulation of the one and two-layer SWQG+1 system that the model captures the asymmetric evolution of vorticity in the one and two-layer shallow model, as well as flow fields with finite divergence.
@@
SWQG+1, like QG+1 in the Boussinesq context, is a PV-based balanced model, where the prognostic variable is PV and all other variables are diagnosed from it. Warn et al. (1995) first considered abstractly the problem of constructing higher-order balanced models and determined that a stable model should not expand the prognostic variable. For SWQG+1 this means PV is not expanded in the small Rossby number. All other variables are expanded and “slaved” to PV by the principle of “PV inversion” of Hoskins et al. (1985). Choosing PV as the central variable is also based on the fact that PV has been crucial in understanding many features of the shallow water model and SWQG (Bretherton & Haidvogel, 1976; Haynes & McIntyre, 1987; Hoskins, 1991; Dritschel & McIntyre, 2008). It is acknowledged that the Boussinesq QG+1 currently lacks an energy conservation principle. The same is the case for the SWQG+1 model.
@@
The rest of the paper is organized as follows. In Section 2, we rederive the one-layer SWQG+1 model using the potentials, which is an alternative form of the model of Warn et al. (1995) and V96. We explore its turbulent free decay, compared to the corresponding shallow water model. It is shown that SWQG+1 captures the emergent negative vorticity skewness in the shallow water model. We also track the energy evolution. It is shown that while theoretical energy conservation law is lacking, the energy of SWQG+1 models the energy evolution of the full shallow water model well. In Section 3, we extend the SWQG+1 model to many layers. A simulation of a baroclinically unstable jet demonstrates SWQG+1 can capture the vorticity asymmetry in two layers. Contrary to the results of the one-layer simulations, the two-layer simulation shows the initial baroclinic growth stage has a cyclonic bias for vorticity. It also shows patterns of vorticity and divergence similar to strain-driven fronts. The vortex stretching mechanism leads to the cyclonic vorticity bias. We conclude in Section 4 and discuss possible applications and extensions to the model.
@@ 2 Single-layer shallow water
This section re-derives the SWQG+1 model, providing an alternative to the Warn et al. (1995)’s approach. By using “potentials” inspired by the Muraki et al. (1999) model for Boussinesq, our form of SWQG+1 involved three elliptic inversions where the elliptic operator to be inverted is the same as the one in SWQG, the Screened Poisson operator. We simulate the free decay of random balanced initial conditions of SWQG+1 model as well as the shallow water model and compare their turbulent statistics. SWQG+1 can capture the vorticity asymmetry in the shallow water model well.
@@ 2.1.1 The shallow water system and its properties
We start with the one-layer shallow water model on a $\beta$-plane and mean layer depth of $H$. The equations for horizontal velocities $(u,v)$ and water layer height perturbation $h$ are
@@
To facilitate future asymptotic analysis but also keep the physical constants for interpretability, we write the equation using the notation of V96. All nondimensional constants are enveloped in curly brackets, and one can ignore them if one wants only the physical equations. One can instead obtain the nondimensional equations by removing all physical constants, here $f$, $\beta$, $g$, and $H$. The nondimensional numbers in this model are the Rossby number:
@@
To prepare for the derivation of the SWQG+1 model, we propose a form of representing the three variables of shallow water using three “potentials”
@@
This procedure and naming are inspired by QG+1 for the Boussinesq model where the three potentials are the vector potential form of the three-dimensional incompressible velocity field (Muraki et al., 1999; Dù et al., 2024). Here, for the shallow water case, the natural counterpart is the two-dimensional Helmholtz decomposition as adopted by Warn et al. (1995) and V96. Though our potentials might look unnatural at first glance, it is useful for the derivation of SWQG+1, where it turns out the elliptic inversion problems for all three potentials are the same as SWQG’s Screened Poisson problems. This uniformity guides the derivation of SWQG+1 in the more complex multi-layer case.
@@ 2.1.2 The shallow water QG and QG+1 model
SWQG+1 is a model that makes the same asymptotic assumption as SWQG but extends it to the next level in Rossby number. That is, we assume the small parameter
@@
$\Phi^{0}$ and $\Phi^{1}$ should have zero mean for all time in the simulation to maintain mass conservation (15c) (assuming without loss of generality that $\left\langle{h}\right\rangle=0$). Therefore for SWQG+1 the $\Phi^{0}$ inversion (18b) should use $q-\left\langle{q}\right\rangle$. Note that with $\beta=0$, the integral of the right-hand side is proportional to the total energy at the QG level after integration by parts
@@
Together, we have the elliptic inversion for $F^{1}$:
@@
In summary, the SWQG+1 inversions provides a way to obtain the physical variables (15) from knowledge of the PV field $q$. It consists of four linear elliptic inversions (19b), (20), (27), and (28). These inversions are all the same Screened Poisson problem. This is typical for a model based on asymptotic expansions. The above derivation of SWQG+1 uses the same ingredient as the model in Warn et al. (1995) and V96, and they are in fact equivalent. The equivalence of the Boussinesq version of the QG+1 model of the model in the Boussinesq version of the V96 model is shown in et al. (2024). We show the equivalency for the shallow water version explicitly in Appendix A. Here, we comment on two points of interest. SWQG+1 is equivalent to the Bolin–Charney balance equation (Bolin, 1955; Charney, 1955).
$f\zeta^{1}-g\nabla^{2}\eta^{1}$ $=-2J(\Phi^{0}_{x},\Phi^{0}_{y})-\left\{\gamma\right\}\beta\Phi_{y}^{0}.$ (29)
This implies SWQG+1 captures the cyclogeotrophic balance. Additionally, we can combine the elliptic inversion for $F^{1}$ and $G^{1}$ to form an elliptic problem for divergence.
$\mathcal{S}(u_{x}+v_{y})$ =$=\left\{\textit{Bu}^{-1}\right\}\frac{f}{gH}\left[J(\Phi^{0},\nabla^{2}\Phi^{0})+\left\{\gamma\right\}\beta\Phi_{x}^{0}\right].$ (30)
This is more commonly called the shallow water $\omega$-equation.
SWQG+1 uses PV as the prognostic variable. Therefore, it retains the PV conservation of (6). However, we have not been able to show that it conserved any definition of energy. Instead, we will demonstrate numerically that the SWQG+1 model generally captures the evolution of the shallow water energy (9), compared to the full model. However, it is occasionally not monotonic for the largest Rossby number explored. Since the simulations use dissipation, this does not prove but indicates that SWQG+1does not conserve the usual shallow water energy. SWQG+1 does not conserve potential enstrophy (7) either, which additionally requires the height to evolve following the height equation (1c). SWQG+1’s height is inverted from PV and does not necessarily follow (1c). We also show from simulations that the potential enstrophy (7) evolved similarly to a full shallow water simulation. Overall, even though SWQG+1 does not explicitly conserve the energy and potential enstrophy, they mirror the behavior of the underlying shallow water model for long-time simulations. We argue that theoretical thinking based on the conserved quantities can be applied to the SWQG+1 simulations.